HUBBLE SPACE TELESCOPE DAILY REPORT #2220 PERIOD COVERED: 0000Z (UTC) 09/30/98 - 0000Z (UTC) 10/01/98 Daily Status Report as of 274/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of WF/PC-2 7712 (WF/PC-2 Cycle 7 Supplemental Darks pt2/3) The WF/PC-2 was used to make a series of Cycle 7 supplemental calibration darks. This dark calibration program obtains three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed Two Sets of STIS/CCD/MA2/WF/PC-2/NIC/1/2/3 8058 (HDF-South Observations) The Space Telescope Imaging Spectrograph (CCD and MA2), the WF/PC-2, and the Near Infrared Camera (Cameras One, Two and Three) were all used to make long exposure Hubble Deep Field (HDFS-22333) observations. This program is designed to provide deep observations of a high-galactic latitude field in the southern continuous viewing zone. The STIS field is centered on a bright QSO at z=2.25. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed WF/PC-2 7620 (WF/PC-2 Cycle 7 Standard Darks) The WF/PC-2 was used to make a series of Cycle 7 standard dark calibration images. This dark calibration program obtains dark frames every week in order to provide data for the ongoing calibration of the CCD dark current rate, and to monitor and characterize the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed STIS/CCD 7948 (CCD Dark and Bias Monitor -- Cont. to Dec. 98) The Space Telescope Imaging Spectrograph (CCD) was used to measure CCD dark current and bias rates for the two supported amplifier settings {GAIN=1 and GAIN=4} in order to gather data for making superdarks and superbiases. The data will be used to track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed FGS 7488 (A Search for Main-Belt Binary Asteroids) The Fine Guidance Sensors were used to make astrometric observations of the main belt asteroid EUNOMIA to search for a possible companion moon/asteroid. The targets was chosen on the basis of an extensive analysis based mainly on peculiar photometric properties suggesting a possible binary nature. The observations were completed as planned, and no problems were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 4 Successful : 4 Scheduled Reacquisitions : 11 Successful : 11 2.2 FHST Updates: Scheduled : 8 Successful : 8 2.3 Operations Notes: The Hubble Deep Field - South (HDF-S) campaign continues. There have been no significant problems reported. The NSSC-1 status buffer was dumped and reset as directed by NS-3, and as expected for HDF-S operations. TTR 20226 was written to document 25 minutes of lost science playback data due to GCMR mismanagement. To recover, the SSR pointers for engineering and science data were reset using ROP SR-3, and SSA transmitter management was accomplished via ROP IC-2. All missing data was received. Using ROP SR-1, the SSR EDAC error counter was reset three times. As a response to an HSTAR filed yesterday, the NSSC-1 memory #3 temperature ground system limit was raised at 273/1945Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH